find diffrence between are of circular table with diameter of 8ft and circular table with diameter 10 ft. round answer to nearst hundredth.
Answer by nerdybill(7384)

Area of a circle = pi(r^2)
circular table w/diameter of 8ft has a radius of 4ft.
circular table w/diameter of 10ft has a radius of 5ft.
area of circular table w/diameter of 8ft = pi(4^2)
area of circular table w/diameter of 10ft = pi(5^2)

Difference in area:
pi(5^2) - pi(4^2)
= pi(25) - pi(16)
= pi(25-16)
= pi(9)
= 3.14(9)
= 28.26